Convert 3.00 x 1020 N2 molecules to mass in grams.
1
Expert's answer
2020-05-24T03:10:04-0400
"n = N\/N_a = 3.00*10^{20}\/6.02*10^{23} = 0.50*10^{-3} mol"

"m = nM = 3.00*10^{-3}*28.00g\/mol = 8.40*10^{-2} g"